OpenStudy (anonymous):

what is the relationship of x and y in 7(x-y)=0

OpenStudy (anonymous):

what is the relationship of x and y in 7(x-y)=0

Directrix (directrix):

This is an odd question. The only thing I know about it is that if 7 and (x-y) multiply to zero, then 7(x-y)=0 Apply Zero Product Property 7 = 0 or x-y =0 @rislam21 If x - y = 0, then solve for x. What do you get? That will tell you the relationship between x and y.

OpenStudy (anonymous):

I find the question strange myself, but how would I solve for x ?

Directrix (directrix):

Picking up from here: 7 = 0 or x-y =0. Observe that 7 cannot equal 0, so forget that. Then, take the equation x - y = 0. Add y to both sides of the equation (Addition Property of Equality). Then, post what you get, and I'll tell you what I got. Okay? @rislam21

OpenStudy (anonymous):

x=y ?

Directrix (directrix):

That is what I got so the relationship of x and y is that they are equal. Do you agree, @rislam21

OpenStudy (anonymous):

Yes, Thank You for the help, I REALLY appreciate it!

Directrix (directrix):

You are welcome.